A USA Today article claims that the proportion of people who believe global warming is a serious issue is 0.78, but given the number of people you've talked to about this same issue, you believe it is greater than 0.78. The hypotheses for this test are Null Hypothesis: p ≤ 0.78, Alternative Hypothesis: p > 0.78. If you randomly sample 28 people and 22 of them believe that global warming is a serious issue, what is your test statistic and p-value?

Question 11 options:

1)

Test Statistic: 0.073, P-Value: 0.942

2)

Test Statistic: 0.073, P-Value: 0.529

3)

Test Statistic: -0.073, P-Value: 0.529

4)

Test Statistic: 0.073, P-Value: 0.471

5)

Test Statistic: -0.073, P-Value: 0.471
Answer #1

Solution :

This is the right tailed test .

The null and alternative hypothesis is

H0 : p = 0.78

Ha : p > 0.78

n = 28

x = 22

\hat p = x / n = 22 / 28 = 0.7875

P0 =  0.78

1 - P0 = 1 - 0.78 = 0.22

Test statistic = z

= \hat p - P0 / [\sqrtP0 * (1 - P0 ) / n]

=0.7875 -  0.78 / [\sqrt(0.78 - 0.22 ) / 28 ]

= 0.073

Test statistic = z = 0.073

P(z > 0.073 ) = 1 - P(z < 0.073 ) = 1 - 0.5291

P-value = 0.471

Option Test Statistic: 0.073, P-Value: 0.471 is correct.